You have been managing a $5 million portfolio that has a beta of 0.85 and a required rate of return of 15.480%. The current risk-free rate is 8%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.15, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

15.48=8+0.85*(market rate-8)

(15.48-8)=0.85*(market rate-8)

market rate=(15.48-8)/0.85+8

=16.8%

Hence required rate for $500,000=8+1.15*(16.8-8)

=18.12%

Hence required rate of portfolio=Respective return*Respective weight

=(5,000,000/5,500,000*15.48)+(500,000/5,500,000*18.12)

=15.72%